OpenStudy (anonymous):

Please help? Simplify 3 ---- x+3 --------- 1 3 --- + --- x+3 x

OpenStudy (anonymous):

In that last question you posted, I submitted a more simplifed solution

OpenStudy (anonymous):

It wouldn't be different for this problem?

OpenStudy (anonymous):

This problem requires a bit of work...

OpenStudy (anonymous):

\[\frac {3} {x+3} \over \frac {1} {x+3} + \frac {3} {x} \]

OpenStudy (anonymous):

\[{ \frac {3} {x+3} \over \frac {1} {x+3} + \frac {3} {x} } = \frac {3} {x+3} \div \frac {1} {x+3} + \frac {3} {x} \]

OpenStudy (anonymous):

The common denominator is (x+3)x

OpenStudy (anonymous):

So multiply each fraction by whatever is missing:

OpenStudy (anonymous):

\[\frac {3} {x+3} \over \frac {1} {x+3} + \frac {3} {x}\] multiply top and bottom by \[x(x+3)\] \[\frac {3} {x+3} \times x(x+3)\over (\frac {1} {x+3} + \frac {3} {x})\times x(x+3)\] \[=\frac{3x}{x+3(x+3)}\]

OpenStudy (anonymous):

for a grand total of \[\frac{3x}{4x+9}\] unless i screwed up the algebra

OpenStudy (anonymous):

I have a question on syntax. In the editor, how do you format the fraction so that it is vertical as opposed to horizontal?

OpenStudy (anonymous):

3x/4x+3 or 3x/4x+9?

OpenStudy (anonymous):

\[{{ \frac {3} {x+3} \over \frac {1} {x+3} + \frac {3} {x} } = \frac {3} {x+3} \div \frac {1} {x+3} + \frac {3} {x} } = \ \frac {3x} {x^2 + 3x} \div \frac {x} {x^2+3x} + \frac {3x+9} {x^2 +3x} \]

OpenStudy (anonymous):

Hero, I think you are missing a couple important parentheses.

OpenStudy (anonymous):

Satellite, you have to teach me how you're able to post multiple lines on the same post

OpenStudy (anonymous):

i will stick with my answer but check it again

OpenStudy (anonymous):

you mean like \[\frac{a}{b}\] \[\frac{c}{d}\]

OpenStudy (anonymous):

Yeah, that...

OpenStudy (anonymous):

i can show you but not here because it will format. i will show you in chAT

OpenStudy (anonymous):

Well yes. A complex fraction.

OpenStudy (anonymous):

\[{{ \frac {3} {x+3} \over \frac {1} {x+3} + \frac {3} {x} } =\frac {3} {x+3} \div (\frac {1} {x+3} + \frac {3} {x} })\]

OpenStudy (anonymous):

i still get \[\frac{3x}{4x+9}\]

OpenStudy (anonymous):

Yeah, satellite has it

OpenStudy (anonymous):

ty i was beginning to worry

OpenStudy (anonymous):

I keep forgetting to multply top and bottom by the lcd

OpenStudy (anonymous):

Yeah I went over it and I think I go with satellite's answer. Thanks guys!

OpenStudy (anonymous):

If you finish it supah fart, you will get the same answer

OpenStudy (anonymous):

Satellites way is most simplest because he begins by multiplying top and bottom by the lcd

OpenStudy (anonymous):

mathsux, I would have gotten there....just would've took a bit longer...I haven't figured out how to post multiple lines yet

OpenStudy (anonymous):

I understand. Thank you hero :)